### v1

# Fine-Tuned Machine Translation Metrics Struggle in Unseen Domains

Abstract
We introduce a new, extensive multidimensional quality metrics (MQM) annotated dataset covering 11 language pairs in the biomedical domain. We use this dataset to investigate whether machine translation (MT) metrics which are fine-tuned on human-generated MT quality judgements are robust to domain shifts between training and inference. We find that fine-tuned metrics exhibit a substantial performance drop in the unseen domain scenario relative to metrics that rely on the surface form, as well as pre-trained metrics which are not fine-tuned on MT quality judgments.
1 Introduction
Automatic metrics are vital for machine translation (MT) research: given the cost and effort required for manual evaluation, automatic metrics are useful for model development and reproducible comparison between research papers (Ma et al., 2019). In recent years, the MT field has been moving away from string-matching metrics like BLEU (Papineni et al., 2002) towards fine-tuned metrics like Comet (Rei et al., 2020), which start with pre-trained models and then fine-tune them on human-generated quality judgments. Fine-tuned metrics have been the best performers in recent WMT Metrics Shared Task evaluations Freitag et al. (2022, 2023) and are recommended by the shared task organizers, who go so far as to say, “Neural fine-tuned metrics are not only better, but also robust to different domains.” (Freitag et al., 2022).
Given the growing popularity of fine-tuned metrics, it is important to better understand their behavior. Here, we examine the question of domain robustness of fine-tuned metrics. Fine-tuned metrics contain extra parameters on top of the pre-trained model which are initialized randomly (or to zero) and then fine-tuned on human-generated MT quality annotations. The primary source of those annotations is from prior WMT metrics shared tasks, and domains in WMT are often carried over from year to year (e.g. news). This raises the question: are fine-tuned metrics in fact robust across any domain (including domains not seen in training)? Or can their apparent strong performance be attributed in part to the artificially good domain match between training and test data?
Figure 1: Automatic machine translation metric performance on the WMT and biomedical domains, averaged across metric types (see Figure 2 for full results).
To answer these questions, we first collect human multidimensional quality metrics (MQM) annotations in the biomedical (bio) domain. Vocabulary overlap and error analysis suggest that this new dataset is distinct from the domains used in WMT. This data covers 11 language pairs and 21 translation systems, with 25k total judgments. In addition to the MQM annotations, we also create new high-quality reference translations for all directions. We release this data publicly, along with code for replication of our experiments.111github.com/amazon-science/bio-mqm-dataset
Next, we examine how different types of metrics perform on our new bio test set relative to the WMT test set. We find that fine-tuned metrics have substantially lower correlation with human judgments in the bio domain, despite other types of metrics having higher correlation in the bio domain (see Figure 1), indicating they struggle with the training/inference domain mismatch. Finally, we present analysis showing that such performance gap persists throughout different stages of the fine-tuning process and is not the result of a deficiency with the pre-trained model.
Architecture Metrics
Surface-Form [Uncaptioned image] BLEU ChrF TER
Pre-trained+Algorithm [Uncaptioned image] BERTScore Prism
Pre-trained+Fine-tuned [Uncaptioned image] Comet UniTE Bleurt
Pre-trained+Prompt [Uncaptioned image] GEMBA AutoMQM
Table 1: Metric types considered in this work. The [Uncaptioned image] components have trainable parameters while [Uncaptioned image] use handcrafted heuristics or algorithms and [Uncaptioned image] decodes from a language model. The ref input omitted in the case of reference-free metrics (i.e. quality estimation).
2 Related Work
Metric types.
Table 1 summarizes the different types of metrics that are commonly used to evaluate MT. The earliest type of MT metrics are Surface-Form metrics, which are purely heuristic and use word- or character-based features. We consider three common Surface-Form metrics: BLEU (Papineni et al., 2002), TER (Snover et al., 2006) and ChrF (Popović, 2015). Metrics like Comet (Rei et al., 2020), Bleurt (Sellam et al., 2020), and UniTE (Wan et al., 2022) start with a pre-trained language model and fine-tune it on human-generated MT quality judgments. We denote these metrics Pre-trained+Fine-tuned. [^2] The WMT metrics task calls these “trained” metrics. Another class of metrics also start with a pre-trained model but do not perform fine-tuning. Examples of such metrics include Prism (Thompson and Post, 2020a, b), which uses the perplexity of a neural paraphraser, and BERTScore (Sun et al., 2022), based on word embedding similarity. We denote such metrics Pre-trained+Algorithem metrics. More recently, metrics like GEMBA (Kocmi and Federmann, 2023) and AutoMQM (Fernandes et al., 2023) have proposed prompting a large language model. We denote these as Pre-trained+Prompt metrics.
Domain specificity.
Domain specificity for MT metrics was first explored by C. de Souza et al. (2014) for Surface-Form metrics. Sharami et al. (2023) brought attention to the issue of domain adaptation for quality estimation (QE), offering solutions with curriculum learning and generating synthetic scores similar to Heo et al. (2021), Baek et al. (2020), and Zouhar et al. (2023). Sun et al. (2022) examined general-purpose natural language generation metrics and documented their bias with respect to social fairness. For word-level QE, Sharami et al. (2023) reported the lack of robustness of neural metrics.
3 New Bio MQM Dataset
We create and release new translations and MQM annotations for the system submissions from 21 participants to the WMT21 bio translation shared task (Yeganova et al., 2021). To explore how different the bio domain is from the WMT22 metric task domains, we computed the vocabulary overlap coefficient between each domain. Bio had the smallest average overlap with the WMT domains (0.436) compared to 0.507, 0.486, 0.507, and 0.582 for e-commerce, news, social, and conversation respectively. See Appendix A for full details and example sentences from each domain.
3.1 Dataset Creation
We created the bio MQM dataset in three steps. Annotations and translations were performed by expert linguists with experience in the medical domain (see Appendix C for full details).
Step 1: Reference re-translation.
The original bio test set consists of bilingual abstracts from crawled academic papers, which might be written by non-native speakers (Névéol et al., 2020) or even MT Thompson et al. (2024). Therefore, we create new professional reference translations.
Step 2: Reference quality.
To ensure a high bar of quality for the reference translations, we ask a separate set of annotators to provide MQM annotations for the new references. Any issues identified by this round of MQM annotation are then fixed by a new set of translators, resulting in the final reference translations that we release in this dataset.
Step 3: MQM annotations.
Finally, we conduct the main MQM annotation on the references and shared task system outputs. In this step, a single annotator rates all translations of a given document (from all systems and the reference). [^3] This allows us to distribute annotation jobs to multiple annotators while still allowing the annotator to access document-level context and ensuring that the whole document is ranked consistently. Our MQM schema follows Freitag et al. (2021) except that we add we add a Critical severity (assigned the same score as Major for backward compatibility). Full annotator instructions are in Appendix D.
WMT Bio
Error severity Critical N/A 8%
Major 26% 44%
Minor 43% 31%
Neutral 31% 16%
Error category Fluency 47% 66%
Accuracy 44% 18%
Terminology 6% 10%
Locale 2% 2%
Other 1% 4%
Error-free segments 45% 72%
Errors per erroneous segment 1.9 2.1
Abs. erroneous segment score -4.1 -7.6
Table 2: Error distribution of our new bio dataset and the existing WMT22 MQM dataset. The MQM annotation scheme for WMT in most cases did not contain the Critical category.
[…]
We compare error distributions on our new bio MQM dataset and the existing WMT MQM dataset in Table 2. Bio MQM contains more Critical/Major errors, and lower absolute scores on average. However, WMT MQM has more overall sentences where an error occurs. Error category distribution also diverges, notably in Fluency and Accuracy.
4 Analysis
Figure 2: Segment-level correlation (Kendall’s $\tau$) between metrics and human judgments on the WMT and bio domain. Pre-trained+Fine-tuned metrics have lower correlation on bio than on WMT, while Surface-Form and Pre-trained+Algorithm tend to have higher correlation. Note that there is no public Comet 22 MQM model.
Here we summarize several experiments aimed at understanding the domain gap observed in Pre-trained+Fine-tuned metrics. Full details are provided in Appendix E and our released code.
Fine-tuned metrics exhibit a performance gap on the unseen domain.
We find (Figure 2) that Pre-trained+Fine-tuned metrics exhibit lower correlation with human judgments on the bio MQM test set than the WMT test set, while both Pre-trained+Algorithem and Surface-Form metrics tend to have higher correlations on bio MQM. This gap suggests that fine-tuned metrics struggle with unseen domains, relative to other metric types. We also observe a very large performance gap for Pre-trained+Prompt metrics. Unfortunately, these metrics rely on closed-source LLMs without published training procedures, so we do not know if the underlying LLMs were trained on WMT metric data.
Domain gap persists throughout the fine-tuning process.
Table 3: Segment-level correlation (Kendall’s $\tau$) between metrics and human judgments on the WMT (top) and bio (bottom) test sets, for Comet with varying epochs of WMT domain DA and MQM training.
Test:WMT MQM epochs
DA epochs
0 1 2 4 8
0 0.118 0.285 0.281 0.279 0.295
1 0.324 0.333 0.318 0.317 0.323
2 0.326 0.337 0.323 0.323 0.325
4 0.322 0.335 0.323 0.322 0.321
8 0.311 0.335 0.324 0.322 0.316
Test:Bio MQM epochs
DA epochs
0 1 2 4 8
0 0.071 0.234 0.229 0.240 0.250
1 0.282 0.280 0.282 0.274 0.270
2 0.270 0.265 0.273 0.268 0.266
4 0.255 0.246 0.258 0.259 0.253
8 0.240 0.242 0.261 0.260
Comet is trained in two steps, first on direct assessment (DA) quality annotations and then on MQM annotations, both from WMT shared tasks. In contrast to catastrophic forgetting Goodfellow et al. (2013); Thompson et al. (2019a, b), where a model starts with good general-domain performance and then overfits while being adapted to a new task or domain, we do not see a sharp dropoff in the bio domain when training on more DA/MQM data. This indicates that the model is a poor bio metric at all stages, as opposed to learning and then forgetting.
Table 3: Segment-level correlation (Kendall’s $\tau$) between metrics and human judgments on the WMT (top) and bio (bottom) test sets, for Comet with varying epochs of WMT domain DA and MQM training.